A rectangular Shape has a 2.50 ± 0.05m length and a 1.4 ± 0.05m width.


i. Determine the Perimeter of this rectangle. State the final answer with its correct uncertainty, as well the correct mumber of significant figures.


ii. Determine the Area of this rectangle. State the final answer with its correct uncertainty, as well the correct number of significant figures. 


iii. If this rectangle had a height of 1.8 +0.02 m, calculate the Volume. State the final answer with its correct uncertainty, as well the correct number of significant

figures.


1
Expert's answer
2021-09-30T18:28:09-0400

i. "P=(2.50+1.40)\plusmn(0.05+0.05)=(3.9\plusmn0.1)\ (m)"


ii. "A=2.5\cdot1.4=3.5\ (m^2)"

"\delta A=3.5\cdot(0.05/2.5+0.05/1.4)=0.2\ (m^2)"


"A=(3.5\plusmn0.2)\ (m^2)"


iii. "V=2.5\cdot1.4\cdot1.8=6.3\ (m^3)"


"\delta V=6.3\cdot(0.05/2.5+0.05/1.4+0.02/1.8)=0.4"


"V=(6.3\plusmn0.4)\ (m^3)"